### v2

# Field Sources for Generalized Ellis-Bronnikov Wormhole

Abstract
This work investigates whether phantom scalar fields combined with nonlinear electrodynamics (NED) can serve as sources for generalized Ellis-Bronnikov wormhole (GEB) spacetimes. This generalization of the Ellis-Bronnikov metric holds significant relevance in the literature and introduces a free parameter $m\geq 2$, recovering the original case when $m=2$. We derive analytical expressions for the scalar field, potential, and NED Lagrangian in magnetic and electric cases. Notably, we explicitly express the Lagrangian as a function of the electromagnetic invariant $\mathcal{F}$ for the magnetic solution. These results provide new insights into the field sources for more general wormhole spacetimes.
I Introduction
In the context of General Relativity (GR), wormholes are exact solutions of Einstein’s equation that describe geometries characterized by the existence of a tunnel capable of connecting two different regions of a spacetime or even two different spacetimes through its throat Morris and Thorne (1988). Although, like black holes, they are spherically symmetric solutions, wormholes are distinguished by being free of event horizons. Moreover, it is known from the literature that the shadows of black holes and horizonless objects, such as naked singularities Bambi and Freese (2009); Ortiz et al. (2015); Shaikh et al. (2018) and non-rotating wormholes Bambi (2013); Azreg-Aïnou (2015), can be similar in some cases Shaikh (2018), which, along with the release of the first black hole shadow image by the Event Horizon Telescope (EHT) Akiyama et al. (2019, 2022), has revitalized interest in studying such objects. In this context, recent works have investigated detection methods for wormholes in both general relativity and alternative theories of gravity, emphasizing their potential astrophysical significance De Falco et al. (2020); De Falco and Capozziello (2023).
It is well-established in the literature that exotic matter, i.e., matter that violates the energy conditions, is required to support such a traversable geometry. In this context, the simplest solution for a traversable Lorentzian wormhole is the Ellis-Bronnikov (EB) wormhole Ellis (1973); Bronnikov (1973), whose metric is given by
$ds^{2}=-dt^{2}+dl^{2}+(l^{2}+b_{0}^{2})d\Omega^{2},$ (1)
where $l$ is known as the proper radial distance, and $b_{0}$ is the radius of the wormhole’s throat. It is generated by an action where gravity is minimally coupled to a free phantom scalar field Bambi and Freese (2009)
$S=\int d^{4}x\left(R-2\epsilon g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right),$ (2)
where $\epsilon=-1$ (phantom) and $\phi=\phi(l)$ is given by
$\phi(l)=\phi_{0}+\arctan\left(\frac{l}{b_{0}}\right).$ (3)
The EB metric is the simplest one describing a traversable wormhole that satisfies the conditions presented in Morris and Thorne (1988). However, in Kar et al. (1995), a generalization of the EB metric is introduced by incorporating a free parameter $m\geq 2$, such that
$ds^{2}=-dt^{2}+dl^{2}+(l^{m}+b_{0}^{m})^{2/m}d\Omega^{2},$ (4)
where $m=2$ recovers the EB wormhole. The following section will discuss Some characteristics of this wormhole family.
On the other hand, the metrics of so-called black bounces Shaikh (2018); Franzin et al. (2021), i.e., spacetimes that transition from regular black holes to wormholes, have been extensively studied recently, with the Simpson-Visser procedure also being applied in other contexts Crispim et al. (2024); Lima et al. (2024, 2023a). An interesting aspect of these geometries is that the field sources generating them are generally composed of nonlinear electrodynamics combined with phantom scalar fields Bronnikov (2024), which has driven the search for field sources of black bounce geometries Bronnikov and Walia (2022); Alencar et al. (2024); Lima et al. (2023b); Bronnikov et al. (2023); Rodrigues and Silva (2023) to become an actively explored topic recently. In the context of wormholes, recent works have presented methods for constructing traversable wormholes supported by nonlinear electrodynamics Cañate and Maldonado-Villamizar (2022), as well as techniques for producing traversable wormholes with electric and magnetic charges without requiring exotic matter Cañate (2024).
Thus, this work investigates whether scalar fields combined with nonlinear electrodynamics can serve as field sources for GEB spacetimes. In the context of NED theories, we employ both magnetic and electric sources, having found analytical expressions for the scalar field, the associated potential, and the Lagrangian of the NED field. In the case of the magnetic solution, we were able to invert the equations to explicitly write the Lagrangian as a function of the invariant $\mathcal{F}=F^{\mu\nu}F_{\mu\nu}$.
This work is structured as follows: In the next section, a brief review of the main characteristics of the GEB metric will be provided; the following section presents our solution, both for the magnetic and electric cases; finally, in the last section, we will offer our discussion and final remarks.
II Generalized Ellis–Bronnikov (GEB) Space-time
As previously mentioned, the Ellis-Bronnikov wormhole metric is a solution to Einstein’s field equation, where a massless scalar field with negative energy acts as the source of curvature. The spherically symmetric metric is typically expressed as
$ds^{2}=-dt^{2}+\frac{dr^{2}}{\left(1-\frac{b(r)}{r}\right)}+r^{2}d\Omega^{2}.$ (5)
Here, $b(r)$ is the shape function that governs the behavior of the wormhole.
The non-zero components of the Einstein tensor for the metric (5) are
$G^{t}_{t}=-\frac{b^{\prime}(r)}{r^{2}},$ (6)
$G^{r}_{r}=-\frac{b(r)}{r^{3}},$ (7)
$G^{\theta}_{\theta}=G^{\varphi}_{\varphi}=\frac{b(r)-rb^{\prime}(r)}{2r^{3}}.$ (8)
For the Ellis-Bronnikov wormhole, the shape function takes on a particular form:
$b(r)=\frac{b_{0}^{2}}{r}$ (9)
where $b_{0}$ is the radius of the wormhole’s throat. The geometry of the solution describes a ”tunnel” that connects two distinct regions of spacetime through the throat. This solution is one of the simplest types of wormhole solutions.
Through a change of coordinates given by
$dl^{2}=\frac{dr^{2}}{\left(1-\frac{b(r)}{r}\right)}$ (10)
it is possible to express the metric as(1).
We can also express the GEB metric in terms of $r$, where the metric takes the form given in (5), where now
$b(r)=r-r^{3-2m}(r^{m}-b_{0}^{m})^{2-\frac{2}{m}}.$ (11)
Setting $t=cte$ and $\theta=\pi/2$ in (4), we obtain the 2D geometry given by
$ds^{2}_{2D}=dl^{2}+(l^{m}+b_{0}^{m})^{2/m}d\varphi^{2}.$ (12)
We can embed this curved 2D geometry into a 3D Euclidean geometry given by the line element
$d\sigma^{2}_{3D}=d\rho^{2}+\rho^{2}d\varphi^{2}+dz^{2}=\left[\left(\frac{d\rho(l)}{dl}\right)^{2}+\left(\frac{dz(l)}{dl}\right)^{2}\right]dl^{2}+\rho^{2}d\varphi^{2},$ (13)
where we identify
$\rho(l)=(b_{0}^{m}+l^{m})^{1/m},$ (14)
$\frac{dz}{dl}=(1-l^{-2+2m}(b_{0}^{m}+l^{m})^{-2+2/m})^{1/2}$ (15)
Numerically integrating the equation (15) with $b_{0}=1$ for different values of $m$ allows us to create the immersion diagram for this GEB, as represented in Fig. 1, where we can observe the classical geometry of a catenoid for $m=2$. For $m>2$, the geometry of the neck of the GEB increasingly tends to become cylindrical.
Figure 1: Embedded diagram for GEB wormhole for $b_{0}=1$ and different values of $m$.
Although the energy conditions continue to be violated for this class of wormholes Kar et al. (1995), the fact that the geometry of the wormhole flattens more rapidly with the increase of $m$ has as one of its consequences the occurrence of resonance in the transmission of massless waves in this geometry, which can be used as a means of obtaining information about the size of the wormhole’s throat Kar et al. (1995), and others studies suggest that this family of wormholes can be used as a template for exploring the gravitational wave physics of exotic compact objects Dutta Roy et al. (2020), and may also serve as a potential black hole mimicker Roy (2022), pending confirmation of stability under all types of perturbations, with initial comparisons of their ringdown characteristics to black holes supporting this idea. Furthermore, it is important to emphasize that this family of solutions has been studied in other contexts beyond GR, such as in braneworld models Sharma and Ghosh (2022a, b, 2021), modified gravity models Godani and Kala (2023); Muniz and Maluf (2022); Nilton et al. (2023), and even in condensed matter contexts de Souza et al. (2022).
In the framework of GR, in the next section, we calculate the fields that serve as sources for this geometry.
III Field Sources for GEB wormhole
III.1 General Relations
Let’s consider an action of the form
$S=\int\sqrt{-g}d^{4}x\left(R-2\epsilon g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+2V(\phi)+\mathcal{L(F)}\right),$ (16)
that is, a scalar field $\phi$ and a nonlinear electromagnetic field minimally coupled with gravity. Furthermore, $\epsilon=-1$ for a phantom scalar field and $\epsilon=+1$ for a canonical scalar field, and $\mathcal{L(F)}$ is the Lagrangian density of the nonlinear electromagnetic field with $\mathcal{F}=F^{\mu\nu}F_{\mu\nu}$, where $F_{\mu\nu}$ is the electromagnetic field tensor.
Varying the action with respect to the metric $g^{\mu\nu}$ yields Einstein’s equation
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=T_{\mu\nu}[\phi]+T_{\mu\nu}[\mathcal{F}],$ (17)
where $T_{\mu\nu}[\phi]$ e $T_{\mu\nu}[\mathcal{F}]$ are, respectively, the energy-momentum tensor associated with the scalar field and the nonlinear electromagnetic field, given by
$T_{\mu\nu}[\phi]=2\epsilon\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\left(\epsilon g^{\lambda\rho}\partial_{\lambda}\phi\partial_{\rho}\phi-V(\phi)\right),$ (18)
$T_{\mu\nu}[\mathcal{F}]=g_{\mu\nu}\frac{\mathcal{L(F)}}{2}-2\mathcal{L_{F}}F^{\alpha}_{\mu}F_{\nu\alpha},$ (19)
where$\mathcal{L_{F}}=d\mathcal{L}/d\mathcal{F}$. Varying the action with respect to the scalar field $\phi$ and the electromagnetic field $F_{\mu\nu}$ we obtain the equations for the fields
$2\epsilon\nabla_{\mu}\nabla^{\mu}\phi=-\frac{dV(\phi)}{d\phi},$ (20)
$\nabla_{\mu}(\mathcal{L_{F}}F^{\mu\nu})=0.$ (21)
III.2 Magnetic Source
In addition to assuming a radial scalar field $\phi=\phi(r)$, we also assume here only the existence of a radial magnetic field, i.e., that the only non-zero components of the electromagnetic tensor $F_{\mu\nu}$ are given by $F_{\theta\varphi}=-F_{\varphi\theta}=q_{m}\sin\theta$, where $q_{m}$ is the charge of the magnetic monopole. With this, we have that the invariant $\mathcal{F}$ is given by
$\mathcal{F}=\frac{2q_{m}^{2}}{r^{4}}.$ (22)
With the metric given in 5, the components of the energy momentum tensors for $\phi$ and $F_{\mu\nu}$ take the form
[…]
$T_{\mu}^{\nu}[\mathcal{F}]=\frac{1}{2}\text{diag}\left(\mathcal{L},\mathcal{L},\mathcal{L}-\frac{4q_{m}^{2}}{r^{4}}\mathcal{L_{F}},\mathcal{L}-\frac{4q_{m}^{2}}{r^{4}}\mathcal{L_{F}}\right).$ (24)
Looking at the above equations, we see that $G^{t}_{t}-G^{r}_{r}=T^{t}_{t}-T^{r}_{r}$ is free of $V$ and $\mathcal{L}$, we have that, using the components of Einstein’s equation given in (6) e (7)
$-2\epsilon\left(1-\frac{b(r)}{r}\right)\phi^{\prime}(r)^{2}=\frac{b(r)}{r^{3}}-\frac{b^{\prime}(r)}{r^{2}}.$ (25)
Using the equation (11), we finally have that the field is given by
$\phi^{\prime}(r)^{2}=\frac{(m-1)}{\epsilon r^{2}\left(1-r^{m}/b_{0}^{m}\right)}.$ (26)
As $r>b_{0}$, for $\phi^{\prime}(r)^{2}>0$, we must have $\epsilon=-1$, that is, a phantom scalar field. Solving the equation (26), we find
$\phi(r)=\phi_{0}+2\frac{\sqrt{m-1}}{m}\arctan\left(\frac{\sqrt{r^{m}-b_{0}^{m}}}{b_{0}^{m/2}}\right),$ (27)
or, in terms of $l$
$\phi(l)=\phi_{0}+2\frac{\sqrt{m-1}}{m}\arctan\left(\frac{l}{b_{0}}\right)^{m/2},$ (28)
where $\phi_{0}$ is a constant that we can fix as zero. From the equation (27), it is simple to see that for $m=2$ we have the scalar field for the Ellis-Bronnikov case Bronnikov (2018). The scalar field (27) is quite similar to the scalar field that serves as the Simpson-Visser spacetime source Bronnikov and Walia (2022), which is also an arc-tangent function, a functional form that is quite recurrent in singularity-free space-times.
Using the equation (20), we can integrate to find the potential $V(r)$, given by
$V(r)=\frac{(m-1)(m-2)b_{0}^{m}}{mr^{2m}}(r^{m}-b_{0}^{m})^{1-2/m}+\frac{2(m-1)(r^{m}-b_{0}^{m})^{1-2/m}}{mb_{0}^{m}}\,_{2}F_{1}\left[2,1-\frac{2}{m},2-\frac{2}{m},1-\frac{r^{m}}{b_{0}^{m}}\right],$ (29)
where ${}_{2}F_{1}(a,b,c,z)$ is the hypergeometric function of $z$. Using the equation (27) to invert, we have that the potential written in terms of the $\phi$ field is given by
$V(\phi)=$ $\frac{(m-1)}{mb_{0}^{2}}\tan^{2-4/m}\left(\frac{m\phi}{2\sqrt{m-1}}\right)\left\{(m-2)\cos^{4}\left(\frac{m\phi}{2\sqrt{m-1}}\right)\right.$
$\left.+2\,\,_{2}F_{1}\left[2,1-\frac{2}{m},2-\frac{2}{m},-\tan^{2}\left(\frac{m\phi}{2\sqrt{m-1}}\right)\right]\right\}$ (30)
In Fig.2 we have plotted the potential as a function of $\phi$ for different values of $m$, where it is possible to observe the periodic character of the potential $V(\phi)$.
Figure 2: Potential $V(\phi)$ for different values of $m$, with $b_{0}=1$.
From Einstein’s equation $(t,t)$ $G^{t}_{t}=T^{t}_{t}$, we have
$-\frac{b^{\prime}(r)}{r^{2}}=V(r)+\left(1-\frac{b(r)}{r}\right)\phi^{\prime}(r)^{2}+\frac{\mathcal{L}(r)}{2},$ (31)
or
$\frac{\mathcal{L}(r)}{2}=-\frac{b^{\prime}(r)}{r^{2}}-V(r)-\left(1-\frac{b(r)}{r}\right)\phi^{\prime}(r)^{2}.$ (32)
Using the expressions for $\phi(r)$ and $V(r)$ found, we obtain the Lagrangian in terms of $r$
$\mathcal{L}(r)=$ $\frac{2}{(r^{m}-b_{0}^{m})^{2/m}}\left\{1-\frac{(m-2)}{m}\left(\frac{b_{0}}{r}\right)^{2m}-\frac{2}{m}\left(\frac{b_{0}}{r}\right)^{m}-\frac{(r^{m}-b_{0}^{m})^{2/m}}{r^{2}}\right.$
$\left.-\frac{2(m-1)(r^{m}-b_{0}^{m})}{mb_{0}^{m}}\,\,_{2}F_{1}\left(2,1-\frac{2}{m},2-\frac{2}{m},1-\frac{r^{m}}{b_{0}^{m}}\right)\right\}$ (33)
In Fig.3 , we have the plot of the Lagrangian as a function of $r$, where we can see a similar behavior for all $m>2$, in which $\mathcal{L}(r)\to$ constant for $r\to\infty$.
Figure 3: Lagrangian $L(r)$ for magnetic case for different values of $m$, with $b_{0}=1$.
We can use the equation (22) to invert the Lagrangian equation and write it in terms of the invariant $\mathcal{F}$, from which we get
$\mathcal{L(F)}=$ $\frac{2}{b_{0}^{2}[(s\mathcal{F}^{-1/4})^{m}-1]^{2/m}}\left\{1-\frac{(m-2)}{m}\left(\frac{\mathcal{F}^{1/2}}{s^{2}}\right)^{m}-\frac{2}{m}\left(\frac{\mathcal{F}^{1/4}}{s}\right)^{m}-\frac{\mathcal{F}^{1/2}}{s^{2}}\left[\left(\frac{s}{\mathcal{F}^{1/4}}\right)^{m}-1\right]^{2/m}\right.$
$\left.+\frac{2(m-1)}{m}\left[1-\left(\frac{s}{\mathcal{F}^{1/4}}\right)^{m}\right]\,\,_{2}F_{1}\left(2,1-\frac{2}{m},2-\frac{2}{m},1-\left(\frac{s}{\mathcal{F}^{1/4}}\right)^{m}\right)\right\},$ (34)
where
$s=\frac{(2q_{m}^{2})^{1/4}}{b_{0}}.$ (35)
As a simplification, we can also define the quantity
$\mathcal{G}=\frac{\mathcal{F}^{1/4}}{s},$ (36)
so that in terms of $\mathcal{G},$ we have
$\mathcal{L(G)}=$ $\frac{2}{b_{0}^{2}(\mathcal{G}^{-m}-1)^{2/m}}\left\{1-\frac{(m-2)}{m}\mathcal{G}^{2m}-\frac{2}{m}\mathcal{G}^{m}-\mathcal{G}^{2}(\mathcal{G}^{-m}-1)^{2/m}\right.$
$\left.+\frac{2(m-1)}{m}(1-\mathcal{G}^{-m})\,\,_{2}F_{1}\left(2,1-\frac{2}{m},2-\frac{2}{m},1-\mathcal{G}^{-m}\right)\right\}.$ (37)
It is easy to see that for $m=2$ we recovered the usual Ellis-Bronnikov case, for which the field source is just a phantom free scalar field. That is, for $m=2$, we have $V(\phi)=1/b_{0}^{2}$ and $\mathcal{L(F)}=-2/b_{0}^{2}$. By substituting this in the action (16) we get
$S(m=2)=\int d^{4}x\left(R-2\epsilon g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+\frac{2}{b_{0}^{2}}-\frac{2}{b_{0}^{2}}\right)=\int d^{4}x\left(R-2\epsilon g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right),$ (38)
that is, for $m=2$, the action reduces to the action of a free scalar field, as expected.
III.3 Electric Source
Alternatively, instead of considering a radial magnetic field generated by a magnetic monopole, we can consider a radial electric field, so that now the non-zero components of the electromagnetic tensor $F_{\mu\nu}$ are given by $F^{tr}=-F^{rt}=E(r)$. With this, the invariant $\mathcal{F}$ is now given by
$\mathcal{F}=-\frac{2E(r)^{2}}{\left(1-\frac{b(r)}{r}\right)}.$ (39)
Also, using the equation (21) we can write the electric field as
$E(r)=\frac{q_{e}}{r^{2}\mathcal{L_{F}}}\sqrt{1-\frac{b(r)}{r}},$ (40)
where $q_{e}$ is the electric charge.
With this, the tensor energy momentum associated with NED is given by
$T[\mathcal{F}]^{\nu}_{\mu}=\frac{1}{2}\text{diag}\left(\mathcal{L}+\frac{4E(r)^{2}\mathcal{L_{F}}}{1-\frac{b(r)}{r}},\mathcal{L}+\frac{4E(r)^{2}\mathcal{L_{F}}}{1-\frac{b(r)}{r}}\mathcal{L},\mathcal{L}\right)$ (41)
Here we proceed analogously to the magnetic case to find the scalar field and the associated potential, which are given by the same expressions (27) and (III.2) as the magnetic case.
Finally, from Einstein’s equation $(\theta,\theta)$ we get
$L(r)=\frac{b(r)-rb^{\prime}(r)}{2r^{3}}-V(r)-\left(1-\frac{b(r)}{r}\right)\phi^{\prime}(r)^{2}.$ (42)
Finally, substituting $\phi(r)$ and $V(r)$, we find that the Lagrangian is given by
$\mathcal{L}(r)=-2\frac{(m-1)(m-2)b_{0}^{m}(r^{m}-b_{0}^{m})^{1-2/m}}{mr^{2m}}-\frac{4(m-1)(r^{m}-b_{0}^{m})^{1-2/m}}{mb_{0}^{m}}\,_{2}F_{1}\left[2,1-\frac{2}{m},2-\frac{2}{m},1-\frac{r^{m}}{b_{0}^{m}}\right].$ (43)
Finally, from the equation $(r,r)$ we can solve for $\mathcal{L}_{F}$, since the electric field in terms of $\mathcal{L}_{F}$ is given in (5). By doing this, we get
$\mathcal{L_{F}}(r)=-\frac{2q_{e}^{2}}{r^{2}[1-r^{2-2m}(r^{m}+(m-2)b_{0}^{m})(r^{m}-b_{0}^{m})^{1-2/m}]}.$ (44)
In Fig.4 we have the plot of the Lagrangian $\mathcal{L}$ as a function of $r$. It is worth noting that, as usually happens in the electrical case, it is not possible to write $\mathcal{L}$ as an explicit function of $\mathcal{F}$, since $\mathcal{F}$ depends on a non-trivial form of $r$. In addition, it is possible to see that, as in the magnetic case, the Lagrangian tends to a constant for asymptotic values of $r$; However, for small values of $r$, $\mathcal{L}(r)$ diverges in the electric case, in contrast to the magnetic case, in which $\mathcal{L}(r)$ tends to a constant to $r\to 0$.
Figure 4: Lagrangian $\mathcal{L}(r)$ for electric case for different values of $m$, with $b_{0}=1$.
Finally, the expression for the electric field is given by
$E(r)=\frac{(r^{m}-b_{0}^{m})^{1-3/m}}{2q_{e}r^{m+1}}\left[r^{4}-b_{0}^{2m}(m-2)r^{4-2m}+b_{0}^{m}(m-3)r^{4-m}-r^{2}(r^{m}-b_{0}^{m})^{2/m}\right].$ (45)
This electric field is obviously not Coulombian. On the limit $r\to\infty$
$E(r)\sim\left(\frac{b_{0}}{r}\right)^{m},$ (46)
which demonstrates that the decay of the $E(r)$ field is faster the higher the value of $m$, which is in agreement with the diagram in Fig.1, which shows that the geometry tends to be flat faster as we increase the value of $m$. In the graph of Fig.5 we have the behavior of the electric field for different values of $m$, where we see that, for $r\to 0$, the electric field does not have a divergent behaviour.
Figure 5: Electric field for different values of $m$, with $b_{0}=1$.
For comparison, in Fig.6 we have plotted the Lagrangian as a function of $r$ for the magnetic case and for the electric case for $m=3$. As previously mentioned, it is possible to see that for asymptotic values of $r$, the Lagrangian tends to the same constant, but assumes completely different behavior for small values of $r$, where the Lagrangian diverges for the electric case.
Figure 6: $\mathcal{L}(r)$ for magnetic and electric case with $b_{0}=1$ and $m=3$.
IV Conclusion
In this work, we have explored whether scalar fields coupled with nonlinear electrodynamics (NED) can serve as field sources for generalized or modified Ellis-Bronnikov (GEB) spacetimes. By employing both magnetic and electric sources in the context of NED theories, we derived analytical expressions for the scalar field, its corresponding potential, and the NED Lagrangian. Notably, for the magnetic solution, we successfully inverted the equations, expressing the Lagrangian as a function of the electromagnetic invariant $\mathcal{F}=F_{\mu\nu}F^{\mu\nu}$. A key observation is that for $m=2$, the solution recovers the well-known Ellis-Bronnikov spacetime sourced by a free phantom scalar field. This connection highlights the versatility of the arc-tangent scalar field, which frequently arises in singularity-free spacetimes, including those of the Simpson-Visser class.
For the electric case, we encountered some differences. Although the Lagrangian also tends to a constant as $r\to\infty$, it diverges for small values of $r$, contrasting with the magnetic case where the Lagrangian remains finite as $r\to 0$. Additionally, the decay of the electric field is influenced by the parameter $m$, with higher values of $m$ leading to a faster approach to flat spacetime, as demonstrated in our results. These distinctions between the magnetic and electric cases offer insight into the behavior of fields in NED-modified spacetimes, emphasizing that the source structure plays a critical role in shaping spacetime geometry.
The ability to express the Lagrangian in terms of the electromagnetic invariant $\mathcal{F}$ for the magnetic case but not for the electric one, due to the dependence of $\mathcal{F}$ on $r$, highlights the richness and complexity of NED as a field source for wormholes. These results not only advance our understanding of the scalar-NED interplay but also open new avenues for further research into different classes of wormhole geometries. Specifically, the approach presented here introduces a novel methodology for seeking field sources in more generalized or modified wormhole spacetimes using NED and scalar fields as core ingredients.
Future investigations can extend this procedure to other geometrical configurations, potentially uncovering new field structures that stabilize exotic spacetimes. The framework developed here serves as a path for exploring a broader range of NED-modified geometries and for the continued search for non-singular, physically viable solutions to wormhole spacetimes.
Acknowledgements
Tiago M. Crispim, G. Alencar and Celio R. Muniz would like to thank Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) for the financial support and Marcos Silva for valuable discussions and insightful comments on the manuscript